OpenStudy (anonymous):

Solve please: W^4-12w^2-2=0 what is the solution w=? type an exact answer using radical as needed use commas to separate answers as needed.

OpenStudy (anonymous):

Let \[y=w^2\] So \[w^4-12w^2-2=0 \Leftrightarrow y^2-12y-2=0\] Solve for y and you'll get two answers. I don't want to type the whole formula here, you should look it up, it's the quadratic formula. The results are \[y=6\pm \sqrt{48}\] But\[w^2=y \Leftrightarrow w=\pm \sqrt{y} \Leftrightarrow w=\pm 6 \pm \sqrt{48}\ or\ w=\pm 6 \mp \sqrt{48}\] So you get 4 roots.

OpenStudy (anonymous):

Sorry I mean \[w=\pm \sqrt{6\pm \sqrt{48}}\ or\ w=\pm \sqrt{6\mp \sqrt{48}}\]

hero (hero):

Jazil, I can help you with you other question...

OpenStudy (anonymous):

ok his answer is not right I need help with this one plse:\[w ^{4}-18w ^{2}-2=0\]

hero (hero):

Oh

hero (hero):

Okay, wait....

OpenStudy (anonymous):

How is my answer not right?

hero (hero):

We'll find out

OpenStudy (anonymous):

it is 38 instead of 48

OpenStudy (anonymous):

hero just work the one i just sent if you can

OpenStudy (anonymous):

OH sorry I made a mistake when I copied it from my calc >.<

hero (hero):

You'll have to come there so we can work on it together